Calculate autonomous consumption expenditure from the following date about an economy which is In equilibrium.
National income = Rs. 1,200
Marginal propensity to save = 0.20
Investment expenditure = Rs. 100 
(Autonomous Consumption Expenditure = 140)

Grade:

2 Answers

Rituraj Tiwari
3 years ago
If MPS=0.20, then
MPC= 1-MPS= 1-0.20= 0.80
Consumption Function is C = c + 0.80 Y where Y in the income in the economy and c= Autonomous consumption.
At equilibrium level of output,
AS=AD
Y= C+I
=> 1,200 = c + 0.80 (1,200) + 100
=> 1,200 = c+ 960 + 100
=> 1,200 = c+ 1,060
=> c = 1,200- 1,060 = Rs.140

thank you
ankit singh
3 years ago

 

If MPS=0.20, then 
MPC= 1-MPS= 1-0.20= 0.80 
Consumption Function is C = c + 0.80 Y where Y in the income in the economy and c= Autonomous consumption. 
At equilibrium level of output,
 AS=AD
Y= C+I 
=> 1,200 = c + 0.80  (1,200)  + 100
=> 1,200 = c+ 960 + 100 
=> 1,200 = c+ 1,060
=> c = 1,200- 1,060 = Rs.140